# Post-Minkowskian Theory Meets the Spinning Effective-One-Body Approach for Bound-Orbit Waveforms

### v1

@@ Abstract
Calibration to numerical-relativity waveforms – As discussed, the accuracy of EOB inspiral-merger-ringdown waveforms can be enhanced through calibration to NR simulations. For the inspiral-plunge stage, this is generally achieved by introducing in the Hamiltonian high-order (still unknown) PN terms, whose coefficients are tuned to NR, and fitting the time of merger (i.e., the (2,2)-mode’s peak time) to NR. In the SEOBNRv5 model Pompili et al. (2023), which was built integrating PN results in the Hamiltonian, three calibration parameters were employed: $(\Delta t_{\rm NR},a_{6},d_{\rm SO})$. The parameter $\Delta t_{\rm NR}$ is defined by $t_{\text{peak }}^{22}=t_{\mathrm{ISCO}}+\Delta t_{\mathrm{NR}}$ (see also Eq. (2)), where $t_{\rm{ISCO}}$ is the time at which $r=r_{\rm{ISCO}}$, with $r_{\rm{ISCO}}$ the radius of the Kerr ISCO Bardeen et al. (1972) with the mass and spin of the remnant BH, as given by NR fitting formula  Jiménez-Forteza et al. (2017); Hofmann et al. (2016). The parameter $a_{6}$ is a 5PN correction to the $A$-potential and $d_{\rm SO}$ is a 4.5PN correction in the gyro-gravitomagnetic coefficients  [^5] More specifically, the SEOBNRv5’s Hamiltonian employs all known 5PN coefficients in the $A$- and $D$-potentials (except for the partially known 5PN local part of the $A$-potential that is fully replaced by the $a_{6}$ calibration coefficient), and the 5.5PN terms in the $Q$-potential. Although the 4.5PN spin-orbit couplings are known, Ref. Pompili et al. (2023) did not use them (but the 3.5PN), because the authors showed that having a spin-orbit calibration parameter at 5.5PN (instead of 4.5PN) is not very effective.. Here, for the SEOBNR-PM model, we do not calibrate high-order PN terms in the non-spinning and spin sectors of the Hamiltonian (3), but we calibrate only the merger’s time through $\Delta t_{\rm NR}$. In future work, we will explore NR calibrations tailored to the particular structure of the PM terms. Henceforth, we compare the PM-informed model with several versions of the most recent PN-GSF–informed SEOBNRv5, with and without calibration.
@@
To calibrate the SEOBNR-PM model, we closely follow the procedure outlined in Refs. Bohé et al. (2017); Pompili et al. (2023); Mihaylov et al. (2023). This procedure essentially consists of determining values of the calibration parameters that minimize a combination of the mismatch and the difference in merger time (defined as the peak of the $(2,2)$-mode amplitude) between EOB and NR waveforms with the same physical parameters ($q,\chi_{1},\chi_{2}$). This is carried out in a Bayesian fashion using the Bilby Ashton et al. (2019) package, and the pySEOBNR code Mihaylov et al. (2023). Finally, we interpolate the best-fit values for each NR simulation across the ($q,\chi_{1},\chi_{2}$) parameter space. As said, in our SEOBNR-PM model, we only calibrate the $\Delta t_{\rm NR}$ parameter (see the Supplemental Material for its expression) using 441 NR simulations of aligned-spin BBHs produced with the pseudo-Spectral Einstein code (SpEC) of the Simulating eXtreme Spacetimes (SXS) Collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017a); Varma et al. (2019a, b); Yoo et al. (2022), which were also employed in Ref. Pompili et al. (2023) for the SEOBNRv5 model. They cover mass ratios $q=m_{1}/m_{2}$ from $1$ to $20$ in the non-spinning limit, and dimensionless spin values going from $-0.998\leq\chi_{i}\leq 0.998$ for $q=1$ to $-0.5\leq\chi_{1}\leq 0.5,~{}\chi_{2}=0$ for $q=15$.
@@
Conclusions – In this Letter, we took advantage of the flexible and efficient pySEOBNR code Mihaylov et al. (2023) and recent prediction for the scattering angle in the EOB formalism Buonanno et al. (2024) to build the first inspiral-merger-ringdown EOB waveform model (SEOBNR-PM) for aligned-spin BHs that uses a PM-informed Hamiltonian (i.e., expanded in $G$, but at all-orders in the velocity). Importantly, we found that the SEOB-PM non-spinning binding energy, computed along an inspiraling trajectory, at 4PM, and its spin-orbit and spin-spin contributions through 5PM, agree remarkably well with the NR data up to 1 GW cycle before merger (see Figs. 2 and  3). The agreement is comparable and in some cases better than SEOBNRv5, which however was calibrated to NR results Pompili et al. (2023). Furthermore, we calibrated SEOBNR-PM to 441 NR simulations provided by the SXS collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017a); Varma et al. (2019a, b); Yoo et al. (2022) by tuning the (2,2)-mode’s peak time (i.e., $\Delta t_{\rm NR}$), and found a median mismatch lower than SEOBNRv5, when the latter is similarly calibrated to NR (see Fig. 4). For now, without optimization, the SEOBNR-PM’s evaluation time is an order of magnitude slower than SEOBNRv5.
@@
Considering the recent attention to the two-body gravitational-scattering problem in quantum-field theory, with a slew of new results produced Bern et al. (2021a, 2022a); Damgaard et al. (2023); Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024), we see the development of the SEOBNR-PM model a watershed moment — the first true application of these methods to an astrophysically relevant inspiral-merger-ringdown waveform model. Yet, this is only a first step. Given the relevant progress at 5PM Driesse et al. (2024), we hope to incorporate the complete 5PM scattering angle into our effective Hamiltonian in the near future. Recent results separating the local from non-local parts of the 4PM angle Dlapa et al. (2024) will likely be crucial for achieving good agreement with NR for highly elliptic bound systems — ultimately, this may be the SEOBNR-PM model’s raison d’être. In light of the progress in PM fluxes Herrmann et al. (2021a, b); Di Vecchia et al. (2021); Heissenberg (2021); Bjerrum-Bohr et al. (2021); Damgaard et al. (2021); Mougiakakos et al. (2021); Manohar et al. (2022); Dlapa et al. (2023b); Jakobsen et al. (2021, 2022c, 2023b), PM corrections could also be fed into the EOB radiation-reaction forces and gravitational modes. The SEOB-PM Hamiltonian and fluxes will also need to be extended to the astrophysically relevant precessing-spin case. We leave these tantalizing prospects for future work.
@@
Acknowledgments – We are grateful to Gustav Uhre Jakobsen, Mohammed Khalil, Jan Plefka and Jan Steinhoff for valuable discussions, and to Raffi Enficiaud for his assistance with scientific computing. The work of G.M. and R.P. was supported by the Deutsche Forschungsgemeinschaft (DFG) Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”.
@@ Appendix A PM-informed effective Hamiltonian in the EOB approach
where ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$ is the inverse tangent integral. This expression also includes $\log(u)$, which is obtained by taking the bound circular-orbit limit (at leading-PN order) of $\log(1-\gamma^{2})$ (i.e., the real part of $\log(\gamma^{2}-1)$). This ensures that the coefficient is real for $\gamma<1$. Finally, to complete the non-spinning dynamics we require the 4PN correction coefficient $\Delta A^{\rm 4PN}$ (5). This includes three numerical coefficients:
@@ Appendix B Interpreting $\gamma$ within $\alpha^{(n)}_{(i,j)}$
A consequence of treating $\gamma$ via Eq. (14) is that it vanishes wherever $H_{\rm Kerr}$ does. For example, in the non-spinning limit, this leads to a pole at the Schwarzschild horizon $r=2M$ in the $\alpha^{(n)}_{(i,j)}$ coefficients (8), which contain powers of $1/\gamma$, and consequently in the $A$-potential. This would not necessarily occur if $\gamma$ was defined as $\gamma=E_{\rm eff}/\mu$, as the effective Hamiltonian does not need to vanish at the horizon. Since the pole appears at a very small separation, it does not prevent evolving the binary until the merger time (which occurs before the time of horizon crossing) and obtaining an accurate inspiral-merger-ringdown waveform. However, the presence of the pole impacts the shape of the $A$-potential (and of the overall Hamiltonian) in the strong-field regime and might limit the flexibility that can be gained from incorporating higher-order calibration parameters in the EOB Hamiltonian. For example, adding a 5PN calibration parameter $a_{6}u^{6}$ to the $A$-potential via Eq. (5), similar to the one used in SEOBNRv5, does not improve the agreement of the non-spinning SEOBNR-PM against NR simulations as effectively as the corresponding parameter in SEOBNRv5. In the future, we will explore different strategies to interpret $\gamma$, and seek for a more suitable choice of the NR calibration parameters tailored to the particular structure of the PM terms, which can enable a more effective NR calibration of the spinning model.

### v2

@@ Abstract
Calibration to numerical-relativity waveforms – As discussed, the accuracy of EOB inspiral-merger-ringdown waveforms can be enhanced through calibration to NR simulations. For the inspiral-plunge stage, this is generally achieved by introducing in the Hamiltonian high-order (still unknown) PN terms, whose coefficients are tuned to NR, and fitting the time of merger (i.e., the (2,2)-mode’s peak time) to NR. In the SEOBNRv5 model Pompili et al. (2023), which was built integrating PN results in the Hamiltonian, three calibration parameters were employed: $(\Delta t_{\rm NR},a_{6},d_{\rm SO})$. The parameter $\Delta t_{\rm NR}$ is defined by $t_{\text{peak }}^{22}=t_{\mathrm{ISCO}}+\Delta t_{\mathrm{NR}}$ (see also Eq. (2)), where $t_{\rm{ISCO}}$ is the time at which $r=r_{\rm{ISCO}}$, with $r_{\rm{ISCO}}$ the radius of the Kerr innermost stable circular orbit (ISCO) Bardeen et al. (1972) with the mass and spin of the remnant BH, as given by NR fitting formula  Jiménez-Forteza et al. (2017); Hofmann et al. (2016). The parameter $a_{6}$ is a 5PN correction to the $A$-potential and $d_{\rm SO}$ is a 4.5PN correction in the gyro-gravitomagnetic coefficients. [^5] More specifically, SEOBNRv5’s Hamiltonian employs all known 5PN coefficients in the $A$- and $D$-potentials (except for the partially known 5PN local part of the $A$-potential that is fully replaced by the $a_{6}$ calibration coefficient), and the 5.5PN terms in the $Q$-potential. Although the 4.5PN spin-orbit couplings are known, Ref. Pompili et al. (2023) did not use them (except for the 3.5PN) because the authors showed that having a spin-orbit calibration parameter at 5.5PN (instead of 4.5PN) is not very effective. Here, for the SEOBNR-PM model, we do not calibrate high-order PN terms in the non-spinning and spin sectors of the Hamiltonian (3), but we calibrate only the merger’s time through $\Delta t_{\rm NR}$. In future work, we will explore NR calibrations tailored to the particular structure of the PM terms. Henceforth, we compare the PM-informed model with several versions of the most recent PN-GSF–informed SEOBNRv5, with and without calibration.
@@
To calibrate the SEOBNR-PM model, we closely follow the procedure outlined in Refs. Bohé et al. (2017); Pompili et al. (2023); Mihaylov et al. (2023). This procedure essentially consists of determining values of the calibration parameters that minimize a combination of the mismatch and the difference in merger time (defined as the peak of the $(2,2)$-mode amplitude) between EOB and NR waveforms with the same physical parameters ($q,\chi_{1},\chi_{2}$). This is carried out in a Bayesian fashion using the Bilby Ashton et al. (2019) package, and the pySEOBNR code Mihaylov et al. (2023). Finally, we interpolate the best-fit values for each NR simulation across the ($q,\chi_{1},\chi_{2}$) parameter space. As said, in our SEOBNR-PM model, we only calibrate the $\Delta t_{\rm NR}$ parameter (see the Supplemental Material for its expression) using 441 NR simulations of aligned-spin BBHs produced with the pseudo-Spectral Einstein code of the Simulating eXtreme Spacetimes (SXS) Collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017b); Varma et al. (2019b, a); Yoo et al. (2022), which were also employed in Ref. Pompili et al. (2023) for the SEOBNRv5 model. They cover mass ratios $q=m_{1}/m_{2}$ from $1$ to $20$ in the non-spinning limit, and dimensionless spin values going from $-0.998\leq\chi_{i}\leq 0.998$ for $q=1$ to $-0.5\leq\chi_{1}\leq 0.5,~{}\chi_{2}=0$ for $q=15$.
@@
Conclusions – In this Letter, we took advantage of the flexible and efficient pySEOBNR code Mihaylov et al. (2023) and recent prediction for the scattering angle in the EOB formalism Buonanno et al. (2024) to build the first inspiral-merger-ringdown EOB waveform model (SEOBNR-PM) for aligned-spin BHs that uses a PM-informed Hamiltonian (i.e., expanded in $G$, but at all orders in the velocity). Importantly, we found that the SEOB-PM non-spinning binding energy, computed along an inspiraling trajectory, at 4PM, and its spin-orbit and spin-spin contributions through 5PM, agree remarkably well with the NR data up to 1 GW cycle before merger (see Figs. 2 and  3). The agreement is comparable and in some cases better than SEOBNRv5, which however was calibrated to NR results Pompili et al. (2023). Furthermore, we calibrated SEOBNR-PM to 441 NR simulations provided by the SXS Collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017b); Varma et al. (2019b, a); Yoo et al. (2022) by tuning the (2,2)-mode’s peak time (i.e., $\Delta t_{\rm NR}$), and found a median mismatch lower than SEOBNRv5, when the latter is similarly calibrated to NR (see Fig. 4). For now, without optimization, the SEOBNR-PM’s evaluation time is an order of magnitude slower than SEOBNRv5.
@@
Considering the recent attention to the two-body gravitational-scattering problem in quantum-field theory, with a slew of new results produced Bern et al. (2021a, 2022a); Damgaard et al. (2023); Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024), we see the development of the SEOBNR-PM model as a watershed moment — the first true application of these methods to an astrophysically relevant inspiral-merger-ringdown waveform model. Yet, this is only a first step. Given the relevant progress at 5PM Driesse et al. (2024), we hope to incorporate the complete 5PM scattering angle into our effective Hamiltonian in the near future. Recent results separating the local from non-local parts of the 4PM angle Dlapa et al. (2024) will likely be crucial for achieving good agreement with NR for highly elliptic bound systems — ultimately, this may be the SEOBNR-PM model’s raison d’être. In light of the progress in PM fluxes Herrmann et al. (2021a, b); Di Vecchia et al. (2021); Heissenberg (2021); Bjerrum-Bohr et al. (2021); Damgaard et al. (2021); Mougiakakos et al. (2021); Manohar et al. (2022); Dlapa et al. (2023b); Jakobsen et al. (2021, 2022c, 2023b), PM corrections could also be fed into the EOB radiation-reaction forces and gravitational modes. The SEOB-PM Hamiltonian and fluxes will also need to be extended to the astrophysically relevant precessing-spin case. We leave these tantalizing prospects for future work.
@@
Acknowledgments – We are grateful to Zvi Bern, Gustav Uhre Jakobsen, Mohammed Khalil, Jan Plefka and Jan Steinhoff for valuable discussions and comments on this Letter, and to Raffi Enficiaud for his assistance with scientific computing. The work of G.M. and R.P. was supported by the Deutsche Forschungsgemeinschaft (DFG) Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”.
@@ Appendix A PM-informed effective Hamiltonian in the EOB approach
where ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$ is the inverse tangent integral. This expression also includes $\log(u)$, which is the logarithmic bound-orbit counterpart to $\log(\gamma^{2}-1)$. The $\log(u)$ can be identified by taking the bound circular-orbit limit (at leading-PN order) of $\log(1-\gamma^{2})$, i.e. the real part of $\log(\gamma^{2}-1)$, ensuring that the coefficient is real for $\gamma<1$. Finally, to complete the non-spinning dynamics we require the 4PN correction coefficient $\Delta A^{\rm 4PN}$ (5). This includes three numerical coefficients:
@@ Appendix B Interpreting $\gamma$ within $\alpha^{(n)}_{(i,j)}$
A consequence of treating $\gamma$ via Eq. (14) is that it vanishes wherever $H_{\rm Kerr}$ does. For example, in the non-spinning limit, this leads to a pole at the Schwarzschild horizon $r=2M$ in the $\alpha^{(n)}_{(i,j)}$ coefficients (8), which contain powers of $1/\gamma$, and consequently in the $A$-potential. This would not necessarily occur if $\gamma$ was defined as $\gamma=E_{\rm eff}/\mu$, as the effective Hamiltonian does not need to vanish at the horizon. Since the pole appears at a very small separation, it does not prevent evolving the binary until the merger time (which occurs before the time of horizon crossing) and obtaining an accurate inspiral-merger-ringdown waveform. However, the presence of the pole impacts the shape of the $A$-potential (and of the overall Hamiltonian) in the strong-field regime and might limit the flexibility that can be gained from incorporating higher-order calibration parameters in the EOB Hamiltonian. For example, adding a 5PN calibration parameter $a_{6}u^{6}$ to the $A$-potential via Eq. (5), similar to the one used in SEOBNRv5, does not improve the agreement of the non-spinning SEOBNR-PM against NR simulations as effectively as the corresponding parameter in SEOBNRv5. We recall that SEOBNRv5 is highly faithful to NR simulations in the non-spinning limit, with a median mismatch of $4.6\times 10^{-5}$ and only $2$ cases having a mismatch $\gtrsim 1\times 10^{-4}$, out of 40 SXS NR waveforms. In the future, we will explore different strategies to interpret $\gamma$, and seek for a more suitable choice of the NR calibration parameters tailored to the particular structure of the PM terms, which can enable a more effective NR calibration of the spinning model.
Figure 5: Mismatch over the binary’s total-mass range $10M_{\odot}\leq M\leq 300M_{\odot}$ for the SEOBNR-PM and SEOBNRv5 models calibrated using only the $\Delta t_{\rm{NR}}$ parameter. The study uses 441 SXS NR waveforms, and focuses on the $(\ell,m)=(2,2)$ mode. Spinning and non-spinning configurations are shown in blue and orange respectively.
@@ Appendix C Fits of the NR-calibration parameters
Appendix D Mismatch in terms of the total mass
In Fig. 5 we show the mismatch over the binary’s total mass range $10M_{\odot}\leq M\leq 300M_{\odot}$ for the SEOBNR-PM and SEOBNRv5 models calibrated using only the $\Delta t_{\rm{NR}}$ parameter, against the 441 SXS NR waveforms used in this work. Spinning and non-spinning configurations are shown in blue and orange, respectively. This provides a complementary picture to Fig. 4, which only presents the maximum mismatch across the total mass range for each configuration.